### v2

@@ Methods
where we can see that the key difference between Equation 8 and Equation 4 is that the latter has a denoising neural network with a dependence on $t$. We will now show that when a masking noise process is used for flow matching, the denoiser’s dependency on $t$ can be dropped, making the two loss functions equivalent.
@@
$x_{t}^{i}=\begin{cases}x_{1}^{i}&\text{if }m_{t}^{i}=0\\ \text{?}&\text{if }m_{t}^{i}=1.\end{cases}$ (9)
@@
$p(x_{1}\mid x_{t},t)$ $=\frac{p(x_{t},t\mid x_{1})p(x_{1})}{\sum_{\tilde{x}_{1}}p(x_{t},t\mid\tilde{x}_{1})p(\tilde{x}_{1})}$ (10)
[…]
$=\frac{\big{[}\prod_{i\in U_{t}}\delta\{x_{t}^{i},x_{1}^{i}\}\big{]}p(x_{1})}{\sum_{\tilde{x}_{1}}\big{[}\prod_{i\in U_{t}}\delta\{x_{t}^{i},\tilde{x}_{1}^{i}\}\big{]}p(\tilde{x}_{1})}$ (12)
$=p(x_{1}\mid x_{t}).$ (13)
@@
where Equation 11 uses the factorization over the dimensions and analyze separately the two cases of whether a position is masked or unmasked, and Equation 12 shows that $p(x_{1}\,|\,x_{t},t)$ does not depend on $t$, giving us Equation 13. Thus, we can rewrite the masking flow matching loss function as:
@@
$\mathcal{L}_{\text{FM}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\,t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p_{\theta}(x_{1}^{i}|x_{t},t)\right]$ (14)
$=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\,t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p_{\theta}(x_{1}^{i}|x_{t})\right]$ (15)
$=\mathcal{L}_{\text{MLM}}.$ (16)
Equivalence between flow matching on discrete state-spaces and order-agnostic autoregressive models
Order-agnostic autoregressive models (OA-AR) differ from standard autoregressive models by training on arbitrary permutations of the sequence ordering [42, 20]. We will now demonstrate that under a masking noise process, the discrete flow matching objective is equivalent to the order-agnostic autoregressive objective.
Let us consider a sequence $x_{1}=(x_{1}^{1},x_{1}^{2},...,x_{1}^{D})$ from our data distribution. In an order-agnostic autoregressive model, we sample a permutation $\sigma$ of indices $\{1,2,...,D\}$ and factorize the joint distribution as:
$p_{\theta}(x_{1})=\prod_{i=1}^{D}p_{\theta}(x_{1}^{\sigma(i)}|x_{1}^{\sigma(<i)})$ (17)
@@
where $x_{1}^{\sigma(<i)}$ denotes $(x_{1}^{\sigma(1)},x_{1}^{\sigma(2)},...,x_{1}^{\sigma(i-1)})$. The OA-AR objective is to maximize the expected log-likelihood over all possible permutations:
@@
$\mathcal{L}_{\text{OA-AR}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\sigma\sim\text{Unif}(\text{Perm}(D))}\left[\sum_{i=1}^{D}\log p_{\theta}(x_{1}^{\sigma(i)}|x_{1}^{\sigma(<i)})\right]$ (18)
@@
$=D\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\sigma\sim\text{Unif}(\text{Perm}(D)),i\sim\text{Unif}(\{1,...,D\})}\left[\log p_{\theta}(x_{1}^{\sigma(i)}|x_{1}^{\sigma(<i)})\right]$ (19)
@@
$=D\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),i\sim\text{Unif}(\{1,...,D\}),\sigma\sim\text{Unif}(\text{Perm}(D))}\left[\log p_{\theta}(x_{1}^{\sigma(i)}|x_{1}^{\sigma(<i)})\right]$ (20)
@@
$=D\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),t\sim\text{Unif}([0,1]),x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log p_{\theta}(x_{1}\mid x_{t})\right]$ (21)
$=D\mathcal{L}_{\text{FM}},$ (22)

### v4

@@ 1.3.1 Equivalence between MFMs and MLMs
where we can see that the key difference between Equation (S7) and Equation (S2) is that the latter has a denoising neural network with a dependence on $t$. We will now show that when a masking noise process is used for flow matching, the denoiser’s dependency on $t$ can be dropped, making the two loss functions equivalent.
@@
$x_{t}^{i}=\begin{cases}x_{1}^{i}&\text{if }m_{t}^{i}=0\\ \text{?}&\text{if }m_{t}^{i}=1.\end{cases}$ (S17)
@@
$p(x_{1}\mid x_{t},t)$ $=\frac{p(x_{t},t\mid x_{1})p(x_{1})}{\sum_{\tilde{x}_{1}}p(x_{t},t\mid\tilde{x}_{1})p(\tilde{x}_{1})}$ (S18)
$=\frac{\cancel{\big[\prod_{i\in M_{t}}(1-t)\prod_{i\in U_{t}}t\big]}\big[\prod_{i\in U_{t}}\delta\{x_{t}^{i},x_{1}^{i}\}\big]p(t)p(x_{1})}{\sum_{\tilde{x}_{1}}\cancel{\big[\prod_{i\in M_{t}}(1-t)\prod_{i\in U_{t}}t\big]}\big[\prod_{i\in U_{t}}\delta\{x_{t}^{i},\tilde{x}_{1}^{i}\}\big]p(t)p(\tilde{x}_{1})}$ (S19)
$=\frac{\big[\prod_{i\in U_{t}}\delta\{x_{t}^{i},x_{1}^{i}\}\big]p(x_{1})}{\sum_{\tilde{x}_{1}}\big[\prod_{i\in U_{t}}\delta\{x_{t}^{i},\tilde{x}_{1}^{i}\}\big]p(\tilde{x}_{1})}$ (S20)
$=p(x_{1}\mid x_{t}).$ (S21)
@@
where Equation (S19) uses the factorization over the dimensions and analyze separately the two cases of whether a position is masked or unmasked, and Equation (S20) shows that $p(x_{1}\,|\,x_{t},t)$ does not depend on $t$, giving us Equation (S21). Thus, we can rewrite the masking flow matching loss function as:
@@
$\mathcal{L}_{\text{FM}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\,t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p_{\theta}(x_{1}^{i}|x_{t},t)\right]$ (S22)
$=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\,t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p_{\theta}(x_{1}^{i}|x_{t})\right]$ (S23)
$=\mathcal{L}_{\text{MLM}}.$ (S24)
1.3.2 Equivalence between MFMs and AO-ARMs
@@
Recall that from Equation (S9) The AO-ARM objective is to maximize the expected log-likelihood over all possible permutations:
@@
$\mathcal{L}_{\text{AO-ARM}}$ $=\mathbb{E}_{x\sim p_{\text{data}}(x),\sigma\sim\text{Unif}(\text{Perm}(D))}\left[\sum_{i=1}^{D}\log p_{\theta}(x^{\sigma(i)}|x^{\sigma(<i)})\right]$
We will show that this objective can be equivalently expressed as a masking discrete flow matching loss:
@@
$\mathcal{L}_{\text{AO-ARM}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\sigma\sim\text{Unif}(\text{Perm}(D))}\left[\sum_{i=1}^{D}\log p_{\theta}(x_{1}^{\sigma(i)}|x_{1}^{\sigma(<i)})\right]$ (S25)
@@
$=D\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\sigma\sim\text{Unif}(\text{Perm}(D)),i\sim\text{Unif}(\{1,...,D\})}\left[\log p_{\theta}(x_{1}^{\sigma(i)}|x_{1}^{\sigma(<i)})\right]$ (S26)
@@
$=D\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),i\sim\text{Unif}(\{1,...,D\}),\sigma\sim\text{Unif}(\text{Perm}(D))}\left[\log p_{\theta}(x_{1}^{\sigma(i)}|x_{1}^{\sigma(<i)})\right]$ (S27)
$=D\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),t\sim\text{Unif}([0,1]),x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log p_{\theta}(x_{1}\mid x_{t})\right]$ (S28)
$=D\mathcal{L}_{\text{FM}},$ (S29)